OpenStudy (friendzone):

what is the probabilty Flipping a Tails AND drawing a King 1/52 15/56 and 1

OpenStudy (johnweldon1993):

Well since 1 has no affect on the other, these are independent events. What is the probability of flipping tails? We have a 1/2 chance What is the probability of drawing a king? Well there are 4 kings in a standard deck of 52 cards... So \[\large \frac{1}{2} \times \frac{4}{52} = ?\]

OpenStudy (friendzone):

@johnweldon1993

OpenStudy (friendzone):

is that possible to dmutilply and get one of those answers